# On the Regularization of Learnable Embeddings for Time Series Processing

### v1

@@ C.2 Forgetting period sensitivity
The introduced forgetting regularization has two hyperparameters: the reset period $k$ and the halting epoch. While the latter can be determined automatically, by monitoring the validation error before each reset, the former should be set empirically. To provide insight on the degree to which the selection of $k$ can impact regularization performance, we perform a sensitivity study on $3$ datasets: METR-LA, PEMS-BAY and AQI. The results are shown in Fig. 5, where we report the mean absolute error (MAE)with varying learning rates and values of $k$. We can see that performance is relatively stable w.r.t. changes in $k$, suggesting that hyperparameter search can be limited to few values (e.g., $3$ for a short period and $30$ for a long one). Another possibility, to avoid searches, is to select the value by hand, after inspecting the learning curves of the unregularized model, in order to select a reset period that allows performance recovery.
@@ Appendix D Detailed experimental setting
All the models were trained with the Adam optimizer (Kingma & Ba, 2015), with a batch size of 64 and up to 300 batches per epoch. We used the Python (Van Rossum & Drake, 2009) programming language, leveraging Torch Spatiotemporal (Cini & Marisca, 2022), Pytorch (Paszke et al., 2019) and Pytorch Lightning (Falcon & The PyTorch Lightning team, 2019) to implement all the experiments. Experiments were scheduled and logged by leveraging Hydra (Yadan, 2019) and Weights and Biases (Biewald, 2020). The mean absolute error (MAE)was used as loss function in all experiments. For the different regularization’s hyperparameters, we heuristically found sensible values (i.e., with stable performance across datasets, models and other hyperparameters) or used those of previous works, in particular: dropout with probability of dropping a connection $p=0.5$, weight of L2 $\lambda_{l2}=0.0001$, weight of L1 $\lambda_{l1}=0.00001$, resetting period of forgetting $k=20$ epochs with $30$ epochs of warm-up, weight of variational regularization $\lambda_{var}=0.00005$ and weight of clustering regularization $\lambda_{clst}=0.0005$. Notice that, in the transfer experiment 3, for variational and clustering regularizations, we used the values adopted in the original paper (Cini et al., 2023b), $\lambda_{var}=0.05$ and $\lambda_{clst}=0.5$ respectively. Any other hyperparameter of variational and clustering regularizations was taken from the original paper as well. Unless specified, we employed and embedding size $d_{e}=32$ for all models. Moreover, all datasets used temporal encodings as additional covariates, in particular, a one-hot encoding of the weekday and sinus and cosinus encodings with daily period.
@@
For the experiments in Tab. 1, optimal hyperparameters for each model, regularization and dataset were found via a grid-search over learning rates values in the range $lr\in[0.00025,0.003]$ and model hidden size value $d_{h}\in[32,256]$, with constraints caused by GPU memory capacity in some settings. Forgetting was halted after $150$ epochs. Each model was trained for 200 epochs and best validation parameters were used for testing. As for the transfer learning experiments in Tab. 3, we followed the settings in Cini et al. (2023b), furthermore, we used a learning rate $lr=0.005$ during training and $lr=0.001$ during fine-tuning. In such scenario, the training lasted up to 150 epochs with 50 epochs of early stopping patience, while fine-tuning lasted up to 1000 epochs with 100 epochs of patience. No temporal encodings were used in the transfer setting. For all the other experiments we used an hidden size $d_{h}=64$ and learning rate $lr=0.00075$, with 300 maximum epochs and early stopping set to 100, while forgetting was halted after 100 epochs. Note that forgetting ignores early stopping while active, to avoid triggering it as a byproduct of the embedding resetting itself.
@@ D.1 Embeddings perturbation
Note that, to avoid penalizing regularizations that are sensible to the weights’ magnitude  (i.e., L1, L2), we consider perturbations that do not change the scale of the learned representations. As this does not hold true for the Noise $\sigma$ perturbation, since it can potentially affect the embeddings magnitude, depending on the standard deviation, we selected most values to be within reason w.r.t. the embeddings magnitude itself. Nonetheless, we also employed a few more extreme values, in order to observe their effects.
@@ Appendix E Effect on simple global models
Table 5: Forecasting error in trasductive setting for MLP model (hidden size 64-128) (5 runs, ±$\pm 1\sigma$). Equal/better than unregularized (+Emb) in bold. Best in red.
@@
Tab. 5 reports the obtained results in case of tuned learning rate and hidden size of $64$ and $128$ units. Compared to models in Tab. 1, the performance of the MLP is worse or comparable at best, however, the regularizations relative effectiveness is similar, w.r.t. each dataset, at hidden size $128$. In the extreme case of the MLP with hidden size $64$, we can observe additional scenarios in which most regularizations are ineffective (i.e., CER-E). Nonetheless, we do not observe catastrophic effects on performance. Notably, with some datasets, i.e., AQI and EngRAD, adding local parameters to the global MLP model hurts performance, while regularization allows an effective exploitation of the embeddings.
Appendix F Additional experiments
@@
In this section, we provide additional experimental results, for completeness w.r.t. what has been shown in the main paper.
@@
Figure 6: Validation curves for different model families (5 runs, ±$\pm 1\sigma$). Curves show the application of dropout regularization at the embedding level or over the entire architecture. Plot names follow the convention model (embedding size) [dataset].
In particular, Fig. 6 complements what has been shown in Fig 2. We can see how, for all the models, regularization at the embedding level seems the preferable choice, even though regularizing the whole model does not impair performance equally in all the settings (e.g., STAtt).
@@
Figure 7: Validation curves for different model families (5 runs, ±$\pm 1\sigma$). Plot names follow the convention model (embedding size) [dataset].
@@
Similarly, Fig. 7 complements the results shown in Fig. 3, with the models that were not shown. We can notice a similar pattern in which dropout (red) can be problematic at smaller embedding sizes (top row). Nonetheless, this seems to be less significant for the STGNN in this specific scenario. In general, the additional figures confirm that dropout (red) and forgetting (blue) lead to different learning curves w.r.t. the un-regularized model (orange), while other regularizations have little impact on this aspect.
@@
Figure 8: Validation curves for different model families (5 runs, ±$\pm 1\sigma$). Plot names follow the convention model (embedding size) [dataset].
@@
Fig. 8 show results obtained by training models with the same setting as Sec. 5.2 on the PEMS-BAY dataset. We can see that, in this case, the validation curves barely plateau. In this scenario, dropout shows potentially problematic effects, similar to those observed for small embedding sizes.
@@
Figure 9: Test performance degradation when adding zero-mean gaussian noise to the embeddings, with increasing variance (5 runs, ±$\pm 1\sigma$, METR-LA). Fixed hidden size and learning rate, $64$ and $0.00075$ respectively.
@@
Figure 10: Test performance degradation on embeddings perturbation (5 runs, ±$\pm 1\sigma$, METR-LA).(Left) random shuffling, (Middle) replaced with their mean, and (Right) replaced by a draw from their sample normal. Fixed hidden size and learning rate, $64$ and $0.00075$ respectively.

### v2

@@ C.2 Forgetting period sensitivity
The introduced forgetting regularization has two hyperparameters: the reset period $k$ and the halting epoch. While the latter can be determined automatically, by monitoring the validation error before each reset, the former should be set empirically. To provide insight on the degree to which the selection of $k$ can impact regularization performance, we perform a sensitivity study on $3$ datasets: METR-LA, PEMS-BAY and AQI. The results are shown in Fig. 5, where we report the mean absolute error (MAE)with varying learning rates and values of $k$. We can see that performance is relatively stable in regard to changes in $k$, suggesting that hyperparameter search can be limited to few values (e.g., $3$ for a short period and $30$ for a long one). Another possibility, to avoid searches, is to select the value by hand, after inspecting the learning curves of the unregularized model, in order to select a reset period that allows performance recovery.
@@ Appendix D Detailed experimental setting
Reproducibility
Python code to reproduce the experiments is available online [^1] https://github.com/LucaButera/TS-embedding-regularization. Datasets will be automatically downloaded or access can be requested to the authors, aside from CER-E and EngRAD. Such datasets can be obtained by contacting the original papers’ authors (EngRAD (Marisca et al., 2024)) or by requesting data access to the appropriate authority (CER-E (CER, 2016)). To specifically reproduce the experiments in Tab 3, regarding the columns for clustering and variational regularizations, access to the code can be requested to the original authors of Cini et al. (2023b). Appendix A contains pointers to request such access. Note that randomized operations in the code are controlled by fixed seeds for reproducibility purposes.
@@
All the models were trained with the Adam optimizer (Kingma & Ba, 2015), with a batch size of 64 and up to 300 batches per epoch. We used the Python (Van Rossum & Drake, 2009) programming language, leveraging Torch Spatiotemporal (Cini & Marisca, 2022), Pytorch (Paszke et al., 2019) and Pytorch Lightning (Falcon & The PyTorch Lightning team, 2019) to implement all the experiments. Experiments were scheduled and logged by leveraging Hydra (Yadan, 2019) and Weights and Biases (Biewald, 2020). The mean absolute error (MAE)was used as loss function in all experiments. For variational and clustering regularizations additional hyperparameters, aside from the regularization strength, i.e., $\lambda_{var}$, $\lambda_{clst}$, were taken from the original paper. Unless specified, we employed and embedding size $d_{e}=32$ for all models. Moreover, all datasets used temporal encodings as additional covariates, in particular, a one-hot encoding of the weekday and sinus and cosinus encodings with daily period.
@@
For the experiments in Tab. 1, optimal hyperparameters, i.e., learning rate $lr$ and hidden size $d_{h}$, for each model and dataset, were found via a grid-search over $lr\in[0.00025,0.00075,0.0015,0.003]$ and $d_{h}\in[32,64,128,256]$, with constraints caused by GPU memory capacity in some settings. Before the grid search for regularized models, regularization’s hyperparameters $\lambda_{l1}$, $\lambda_{l2}$, $\lambda_{var}$ and $\lambda_{clst}$, were set according to the following procedure. For each model and dataset (we here consider three datasets: METR-LA, PEMS-BAY and AQI), the validation error is computed for regularization strength values in $\left[0.01,0.001,0.0001,0.00001\right]$, while keeping the $lr$ and $d_{h}$ fixed to the optimal values for the respective un-regularized model. For each regularization hyperparameter, the corresponding validation results are ranked and the average rank is computed across models and datasets. The parameter with the top rank is selected and kept fixed. This resulted in: weight of L2 $\lambda_{l2}=0.0001$, weight of L1 $\lambda_{l1}=0.00001$, weight of variational regularization $\lambda_{var}=0.00005$ and weight of clustering regularization $\lambda_{clst}=0.0005$. Dropout’s probability of dropping a connection was set to $p=0.5$, which is a middle ground between soft and aggressive dropout. Moreover, we set the resetting period of forgetting to $k=20$ epochs with $30$ epochs of warm-up. The analysis in Sec. C.2 shows how the method is not particularly sensitive in this regard. Forgetting was halted after $150$ epochs. Each model was trained for $200$ epochs and the best validation parameters were used for testing.
For the experiments in Tab. 3, we used the STGNN in Eq. 10 and set $d_{h}=64$, $\lambda_{var}=0.05$ and $\lambda_{clst}=0.5$ as in Cini et al. (2023b), while for the other regularizations we used the same hyperparameters as in Tab. 1. Furthermore, we used a learning rate $lr=0.005$ during training and $lr=0.001$ during fine-tuning. The training lasted up to $150$ epochs with $50$ epochs of early stopping patience, while fine-tuning lasted up to $1000$ epochs with $100$ epochs of patience. No temporal encodings were used in this setting.
For experiments in Sec. 5.2 and Sec. 5.4 we used set $d_{h}=64$ and $lr=0.00075$. Regularization’s hyperparameters were set to the values found for Tab. 1. Training lasted up to $300$ epochs, early stopping patience was set to $100$, while forgetting was halted after $100$ epochs. Note that the forgetting regularization ignores early stopping for the whole duration before being halted; this is to avoid triggering early stopping as a byproduct of the parameter resetting itself.
@@ D.1 Embeddings perturbation
Note that, to avoid penalizing regularizations that are sensible to the weights’ magnitude  (i.e., L1, L2), we consider perturbations that do not change the scale of the learned representations. As this does not hold true for the Noise $\sigma$ perturbation, since it can potentially affect the embeddings magnitude, depending on the standard deviation, we selected most values to be within reason relative to the embeddings magnitude itself. Nonetheless, we also employed a few more extreme values, in order to observe their effects.
@@ Appendix E Effect on simple global models
Table 5: Forecasting error in transductive setting for MLP model (hidden size 64-128) (5 runs, ±$\pm 1std$). Equal to/better than unregularized (+ Emb) in bold. Best in red. ‘+ Regularization’ denotes the addition of that specific regularization only, on top of ‘+ Emb.’.
@@
Tab. 5 reports the obtained results in case of tuned learning rate and hidden size of $64$ and $128$ units. Compared to models in Tab. 1, the performance of the MLP is worse or comparable at best, however, the regularizations relative effectiveness is similar, considering each dataset, at hidden size $128$. In the extreme case of the MLP with hidden size $64$, we can observe additional scenarios in which most regularizations are ineffective (i.e., CER-E). Nonetheless, we do not observe catastrophic effects on performance. Notably, with some datasets, i.e., AQI and EngRAD, adding local parameters to the global MLP model hurts performance, while regularization allows an effective exploitation of the embeddings.
Appendix F Effect of combining different regularizations
To further investigate whether different regularization techniques can be combined for improved performance, we applied L2 and dropout on top of the approaches specifically designed for the regularization of local embeddings, i.e., clustering, variational and forgetting regularizations. This covers the most reasonable combinations, as L2 and dropout are commonly used in combination with other regularization approaches. In particular, we consider the same transductive setting as in Sec. 5.1, and adopt the optimal hyperparameter configuration (see Appendix D for details) previously found for clustering, variational, and forgetting, respectively.
Table 6: Forecasting test error under optimal model size and learning rate from Tab. 1 (STGNN, 5 runs, ±$\pm 1std$) when adding additional regularizations. Combinations equal to or better than the corresponding base regularization are in bold. The best-performing method within each dataset and base regularization is in red. Regularization methods are always considered applied on top of the ‘+ Emb.’ model.
Dataset METR-LA PEMS-BAY AQI
Model MAE $\downarrow$ MAE $\downarrow$ MAE $\downarrow$
+ Clust. $3.025_{\pm.012}$ $1.580_{\pm.005}$ $11.876_{\pm.053}$
+ Clust. + L2 ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{3.019}}_{\pm 0.008}$ $1.581_{\pm 0.008}$ $\mathbf{11.767}_{\pm 0.023}$
+ Clust. + Drop. $3.052_{\pm 0.007}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{1.572}}_{\pm 0.009}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{11.712}}_{\pm 0.037}$
+ Vari. $\mathbf{3.013}_{\pm.005}$ $\mathbf{1.566}_{\pm.003}$ $11.768_{\pm.026}$
+ Vari. + L2 $3.037_{\pm 0.009}$ $1.581_{\pm 0.007}$ $11.789_{\pm 0.022}$
+ Vari. + Drop. $3.047_{\pm 0.008}$ $1.576_{\pm 0.004}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{11.675}}_{\pm 0.023}$
+ Forg. $3.050_{\pm.017}$ $1.578_{\pm.006}$ $11.793_{\pm.040}$
+ Forg. + L2 $\mathbf{3.045}_{\pm 0.012}$ $\mathbf{1.576}_{\pm 0.007}$ $\mathbf{11.764}_{\pm 0.023}$
+ Forg. + Drop. $3.087_{\pm 0.005}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{1.570}}_{\pm 0.003}$ ${\color[rgb]{0.8078,0.2745,0.2745}\mathbf{11.755}}_{\pm 0.028}$
Tab. 6 shows the obtained results. We can see that combining different regularization techniques can be beneficial. Noticeably, combining variational regularization and dropout yielded a new overall best score on the AQI dataset. However, improvements appear mostly case-dependent, without emerging clear patterns.
Appendix G Additional experiments
@@
For completeness, in this section, we provide experimental results in addition to what has been shown in the main paper.
@@
Figure 6: Validation curves for different model families (5 runs, ±$\pm 1std$). Curves show the application of dropout regularization at the embedding level or over the entire architecture. Plot names follow the convention model (embedding size) [dataset].
Fig. 2 shows an example of how regularizing local parameters only can have a different impact than regularizing all parameters on the training of an STGNN model. Fig. 6 complements what has been shown in such an example by providing results for the other two models, i.e., RNN and STAtt. A similar pattern emerges for all the considered models, though with different degrees.
@@
Figure 7: Validation curves for different model families (5 runs, ±$\pm 1std$). Plot names follow the convention model (embedding size) [dataset].
@@
Similarly, Fig. 7 complements the results shown in Fig. 3, with the models that were not shown. We can notice a similar pattern in which dropout (red) can be problematic at smaller embedding sizes (top row). Nonetheless, this seems to be less significant for the STGNN in this specific scenario. In general, the additional plots confirm that dropout (red) and forgetting (blue) lead to different learning curves compared to the un-regularized model (orange), while other regularizations have little impact on this aspect.
@@
Figure 8: Validation curves for different model families (5 runs, ±$\pm 1std$). Plot names follow the convention model (embedding size) [dataset].
@@
Fig. 8 shows results obtained by training models with the same setting as Sec. 5.2 on the PEMS-BAY dataset. We can see that, in this case, the validation curves barely plateau. In this scenario, dropout shows potentially problematic effects, similar to those observed for small embedding sizes.
@@
Figure 9: Test performance degradation when adding zero-mean Gaussian noise to the embeddings, with increasing variance (5 runs, ±$\pm 1std$, METR-LA). Fixed hidden size and learning rate, $64$ and $0.00075$ respectively.
@@
Figure 10: Test performance degradation on embeddings perturbation (5 runs, ±$\pm 1std$, METR-LA).(Left) random shuffling, (Middle) replaced with their mean, and (Right) replaced by a draw from their sample normal. Fixed hidden size and learning rate, $64$ and $0.00075$ respectively.